# TCIP: Threshold-Controlled Iterative Pyramid Network for Deformable Medical Image Registration

### v1

@@ Abstract
Although pyramid networks have demonstrated superior performance in deformable medical image registration, their decoder architectures are inherently prone to propagating and accumulating anatomical structure misalignments. Moreover, most existing models do not adaptively determine the number of iterations for optimization under varying deformation requirements across images, resulting in either premature termination or excessive iterations that degrades registration accuracy. To effectively mitigate the accumulation of anatomical misalignments, we propose the Feature-Enhanced Residual Module (FERM) as the core component of each decoding layer in the pyramid network. FERM comprises three sequential blocks that extract anatomical semantic features, learn to suppress irrelevant features, and estimate the final deformation field, respectively. To adaptively determine the number of iterations for varying images, we propose the dual-stage Threshold-Controlled Iterative (TCI) strategy. In the first stage, TCI assesses registration stability and with asserted stability, it continues with the second stage to evaluate convergence. We coin the model that integrates FERM and TCI as Threshold-Controlled Iterative Pyramid (TCIP). Extensive experiments on three public brain MRI datasets and one abdomen CT dataset demonstrate that TCIP outperforms the state-of-the-art (SOTA) registration networks in terms of accuracy, while maintaining comparable inference speed and a compact model parameter size. Finally, we assess the generalizability of FERM and TCI by integrating them with existing registration networks and further conduct ablation studies to validate the effectiveness of these two proposed methods.
@@ 1 Introduction
Deformable image registration aims to align the corresponding anatomical structures between a fixed and moving image pair by estimating a dense non-linear deformation field, with widespread real-world applications in medical image analysis (Cai et al. 2020; Rana, Hampson, and Dobie 2019). To estimate the deformation field, conventional methods perform separate optimizations for each image pair, incurring high computational costs (Fischer and Modersitzki 2008).
@@
To improve efficiency, prior studies have leveraged Deep Learning (DL) methods for deformable image registration (Jia et al. 2022, 2023), with pioneering studies (Balakrishnan et al. 2018, 2019) primarily employing the U-Net networks to estimate the deformation field and warp the moving image. However, U-Net networks typically produce a single deformation field at the output layer, limiting their ability to accurately capture deformations that are large in both magnitude and complexity (Lv et al. 2022). To better address this challenge, certain studies adopted a cascaded registration scheme involving multiple subnetworks, each operating at a specific resolution (i.e., scale) to extract features and estimate a corresponding deformation field (De Vos et al. 2019; Mok and Chung 2020). This approach, while effective in handling large deformations, often redundantly re-extracts features at each subnetwork and fails to leverage previously encoded features (Wang, Ni, and Wang 2024).
@@
To minimize unnecessary feature extraction, recent studies have proposed pyramid models (Kang et al. 2022; Liu et al. 2022; Ma et al. 2024; Meng et al. 2024) that employ a single encoder to extract multi-scale feature maps. Subsequently, at each scale, the decoding layer fuses the feature maps extracted from both the fixed and moving image branches to estimate the deformation field. The estimated deformation field is then passed to the subsequent decoding layer (see Figure 1(a)), thereby achieving coarse‑to‑fine estimation. Although pyramid architectures facilitate multi-scale feature extraction, their decoders may exacerbate the propagation and accumulation of anatomical structure misalignments, especially those stemmed from irrelevant information in deep, coarse‑scale features. This exacerbation ultimately constrains the overall registration performance (Zhou et al. 2023). In addition, existing pyramid models often adopt vanilla residual decoders, which often lead to unnecessary computation resource utilization. Moreover, in scenarios involving large deformations, a one-pass estimation per scale typically fails to yield an accurate deformation field, necessitating iterative estimations at each scale. Consequently, SOTA models perform multiple estimations at each scale to iteratively optimize the deformation fields. For example, Wang, Ni, and Wang (2024) adopted a fixed number of iterations, while Ma et al. (2024) introduced a single-stage mechanism to dynamically determine the number of iterations at each scale. However, due to the diversity among image pairs, neither a fixed number of iterations nor a single-stage criterion may effectively adapt to varying deformation requirements, often resulting in premature termination or excessive iterations that degrades registration performance (see Tables 1 and 2).
@@
Figure 1: Illustration of the differences in decoding processes between existing pyramid models and our proposed TCIP. Sub-figure (a) portrays the existing models (Wang, Ni, and Wang 2024) and (Ma et al. 2024), which neither effectively prevent propagation and accumulation of anatomical structure misalignments nor adaptively determine the number of iterations. Sub-figure (b) depicts the proposed TCIP. At each decoding layer, TCIP employs our designed FERM to learn to suppress irrelevant features, thereby reducing misalignment accumulation. Subsequently, TCIP employs our novel dual-stage TCI strategy to adaptively determine the number of iterations for guiding FERM in the progressive optimization process of the deformation field.
@@
To mitigate misalignment propagation and accumulation in pyramid architectures, we propose the Feature-Enhanced Residual Module (FERM) as the core component of the decoding layer, which can emphasize informative features. Specifically, FERM comprises three sequential blocks (see Figure 1(b)), namely Feature Fusion Block (FFB), Squeeze Excitation Block (SEB), and Deformation Field Estimator (DeF). Initially, at the $l$th scale, the residual FFB extracts anatomical semantic features from both the fixed and moving images, while preserving detailed original information. Subsequently, SEB applies channel-wise attention to reweight these extracted features (see Section 3.2), learning to suppress irrelevant ones. Finally, DeF estimates the deformation fields. By adopting this design, FERM achieves high-level registration performance with fewer parameters comparing to a vanilla residual decoder. Moreover, to adaptively determine the number of iterations for diverse image pairs, we propose the dual-stage Threshold-Controlled Iterative (TCI) strategy, which employs a sliding window to dynamically track the estimated deformation fields and quantifies each by the similarity difference between the warped image $I_{m}^{w}$ and the fixed image $I_{f}$. TCI then assesses the stability of deformation (i.e., first stage) by computing the standard deviation of all similarity differences within this window. Upon reaching stability (see Section 3.3), TCI further evaluates convergence (i.e., second stage) by measuring the change between similarity differences of the two most recently estimated deformation fields within this window. TCI is designed to be efficient to enable adaptive determination of the number of iterations for each scale of the underlying image pair, without significantly increasing registration time (see Section 4.1). We coin the network that integrates FERM and TCI as the Threshold-Controlled Iterative Pyramid (TCIP) network.
@@
i) We propose the lightweight FERM as the core component of the decoding layer, which integrates residual blocks and channel-wise attention to extract anatomical features and learn to suppress irrelevant information, thereby mitigating misalignment accumulation.
ii) We propose the dual-stage TCI strategy, which monitors the stability and convergence of recent registration results to adaptively determine the number of iterations for image pairs with varying deformation requirements.
iii) To assess the performance of TCIP, we conduct extensive experiments on three public brain Magnetic Resonance Imaging (MRI) datasets and one abdomen Computed Tomography (CT) dataset. The results demonstrate that TCIP outperforms SOTA registration models in accuracy across all datasets, while achieving on-par inference speed and a compact model size. Moreover, generalization and ablation studies further validate the effectiveness of FERM and TCI.
@@ 2 Related Work
Deep Learning (DL) Registration Models: DL models have become the mainstream for registration tasks due to their advantages in inference speed and accuracy compared to conventional registration methods (Balakrishnan et al. 2019). Pioneer DL models primarily employed U-Net architectures, which estimate only a single deformation field at the final output layer and consequently struggle to accommodate complex deformations (Wang and Wang 2025). To better overcome this limitation, certain studies (De Vos et al. 2019; Zhao et al. 2019a, b) proposed cascaded registration networks, which employ multiple sub-networks to extract features at distinct resolutions and estimate the corresponding deformation fields. Although these models achieve moderate accuracy improvements, they perform repeated encoding and decoding, leading to redundant feature extraction and elevated computing overhead (Ma et al. 2024).
@@
Figure 2: Overall TCIP model architecture. A weight-sharing encoder first extracts multi-scale feature maps $\{F_{l}\}$ and $\{M_{l}\}$ for the fixed image $I_{f}$ and the moving image $I_{m}$, respectively. In the subsequent decoding layers, TCIP employs FERM to emphasize informative features and estimate deformation field $\phi_{l}$, while TCI adaptively determines the number of iterations for each decoding layer, enabling progressive optimization of $\phi_{l}$. Notably, the encoder processes higher-resolution feature maps first, while the decoder begins deformation estimation on the lowest‑resolution maps and proceeds to finer scales.
@@
To minimize redundant feature processing, recent studies (Liu et al. 2022) proposed the pyramid network, which adopts a single encoder followed by multiple decoding layers to estimate hierarchical deformation fields for continuous coarse‑to‑fine warping images. Although this architecture facilitates effective multi-scale feature extraction and progressive deformation estimation, it may exacerbate the propagation and accumulation of anatomical structure misalignments, especially those stemmed from irrelevant information in deep, coarse‑scale features which ultimately affect the overall registration performance (Zhou et al. 2023). To better overcome this limitation, we propose FERM, which learns to suppress irrelevant features through feature weighting in the decoding layers, thereby mitigating misalignment propagation. Moreover, FERM excels at capturing anatomical details, ensuring the accuracy and consistency of deformation fields at each scale (see Section 3.2).
@@
Iterative Strategy in Pyramid Models: For small deformations, prior pyramid models (Kang et al. 2022) estimated only a single deformation field per decoding layer (each corresponding to one scale). Consequently, each scale requires iterative estimations to refine the deformation field. For example, Wang, Ni, and Wang (2024) proposed RDP, which performs a predefined number of iterations for each scale to refine the deformation field. Inevitably, this fixed number of iterations lacks adaptability to varying deformation requirements. To adaptively determine the number of iterations, Ma et al. (2024) proposed a single-stage mechanism, which evaluates convergence by monitoring differences between two successive registration results. However, relying on a single-metric check only captures partial information and may lead to premature iteration termination, potentially compromising the overall registration accuracy. To dynamically determine when to stop iterative optimization, we propose a dual-stage mechanism, which evaluates the stability of historical registration results and convergence subsequently (see Section 3.3).
3 Threshold-Controlled Iterative Pyramid
@@
The architecture of TCIP is schematically illustrated in Figure 2. In this section, we first introduce FERM, which excels at extracting anatomical structure details and learn to suppress irrelevant features to mitigate misalignment. We then introduce the dual-stage TCI strategy, which adaptively determines the number of iterations for optimization.
3.1 Overview of TCIP
@@
The proposed TCIP comprises a multi-scale encoder and a multi-layer decoder. TCIP first employs a weight-sharing feature encoder to extract multi-scale feature map sets $\{F_{l}\}$ and $\{M_{l}\}$, where $l\in\{1,...,4\}$ (following prior arts in this field: (Kang et al. 2022; Ma et al. 2024)), from the fixed image $I_{f}$ and moving image $I_{m}$, respectively. Specifically, TCIP adopts the encoder used in (Ma et al. 2024), which comprises four convolutional blocks, each comprising a 3D convolutional layer with a LeakyReLU activation function, followed by average pooling that halves the spatial dimensions of the feature maps. Then, for the $l$th scale (i.e., the $l$th decoding layer), TCIP applies FERM to process $\{F_{l}\}$ and $\{M_{l}\}$, extracting anatomical structure details and learning to suppress irrelevant features, to estimate the corresponding deformation fields (see Section 3.2). To adaptively determine the number of iterations in each decoding layer, TCIP relies on TCI, which monitors the similarity difference of the recently estimated registration results within a sliding window (see Section 3.3).
Figure 3: Depiction of the proposed FERM, comprising the FFB, SEB, and DeF modules. FFB is responsible for extracting anatomical structure details. SEB suppresses irrelevant features to mitigate anatomical structure misalignment accumulation. DeF estimates the deformation fields.

### v2

# iPEAR: Iterative Pyramid Estimation with Attention and Residuals for Deformable Medical Image Registration

@@ Abstract
Existing pyramid registration networks may accumulate anatomical misalignments and lack an effective mechanism to dynamically determine the number of optimization iterations under varying deformation requirements across images, leading to degraded performance. To solve these limitations, we propose iPEAR. Specifically, iPEAR adopts our proposed Fused Attention-Residual Module (FARM) for decoding, which comprises an attention pathway and a residual pathway to alleviate the accumulation of anatomical misalignment. We further propose a dual-stage Threshold-Controlled Iterative (TCI) strategy that adaptively determines the number of optimization iterations for varying images by evaluating registration stability and convergence. Extensive experiments on three public brain MRI datasets and one public abdomen CT dataset show that iPEAR outperforms state-of-the-art (SOTA) registration networks in terms of accuracy, while achieving on-par inference speed and model parameter size. Generalization and ablation studies further validate the effectiveness of the proposed FARM and TCI.
Machine Learning, ICML
@@ 1 Introduction
In single-modality settings, deformable image registration aims to align the corresponding anatomical structures between a fixed and moving image pair within the same imaging modality by estimating a dense non-linear deformation field, with widespread real-world applications in medical image analysis (Cai et al., 2020; Rana et al., 2019). To estimate the deformation field, conventional methods perform separate optimizations for each image pair, incurring high computational costs (Fischer and Modersitzki, 2008).
@@
To improve efficiency, deep learning methods have been widely adopted for deformable image registration, with pioneering works (Balakrishnan et al., 2018; De Vos et al., 2019) mainly adopting U-Net and Cascaded frameworks to estimate deformation fields and warp moving images. U-Net networks typically predict a single deformation field at the output layer, which limits their ability to model large and complex deformations (Lv et al., 2022). Cascaded schemes address this issue by employing multiple subnetworks at different scales to progressively refine the deformation field, but they often re-extract features redundantly and underutilize previously encoded representations (Wang et al., 2024).
@@
To minimize unnecessary feature extraction, recent studies have proposed pyramid models (Kang et al., 2022; Liu et al., 2022) that employ a single encoder to extract multi-scale feature maps. Subsequently, at each scale, the decoding layer fuses the feature maps extracted from both the fixed and moving image branches to estimate the deformation field. The estimated deformation field is then passed to the subsequent decoding layer (see Figure 1(a)), thereby achieving coarse‑to‑fine estimation. Although pyramid architectures facilitate multi-scale feature extraction, their decoders may exacerbate the propagation and accumulation of anatomical structure misalignments, especially those stemmed from irrelevant information in deep, coarse‑scale features. This exacerbation ultimately constrains the overall registration performance (Zhou et al., 2023). In addition, existing single-modality pyramid models often adopt vanilla residual decoders, which limit their performance (see Tables 1 and 2). Moreover, in scenarios involving large deformations, a one-pass estimation per scale typically fails to yield an accurate deformation field, necessitating iterative estimations at each scale. Consequently, SOTA models perform multiple estimations at each scale to iteratively optimize the deformation fields. For example, Wang et al. (2024) adopted a fixed number of iterations, while Ma et al. (2024) introduced a single-stage mechanism to dynamically determine the number of iterations at each scale. However, due to the diversity among image pairs, neither a fixed number of iterations nor a single-stage criterion may effectively adapt to varying deformation requirements, often resulting in premature termination or excessive iterations that degrade registration performance (see Section 4.1).
@@
Figure 1: Illustration of the differences in decoding processes between existing pyramid models and our proposed iPEAR. Sub-figure (a) portrays the existing models (Wang et al., 2024) (left) and (Ma et al., 2024) (right), which do not specifically prevent propagation and accumulation of anatomical structure misalignments nor adaptively determine the number of iterations. Sub-figure (b) depicts our proposed iPEAR. At each decoding layer, iPEAR employs our designed FARM to learn to suppress irrelevant features, thereby reducing misalignment accumulation. Subsequently, iPEAR employs our novel dual-stage TCI strategy to adaptively determine the number of iterations for guiding FARM in the progressive optimization process of the deformation field.
@@
To mitigate misalignment propagation and accumulation in pyramid architectures, we propose the Fused Attention-Residual Module (FARM) as the core component of the decoding layer to emphasize informative features. Specifically, FARM follows a dual-pathway design, including an Attention Pathway (AP) and a Residual Pathway (RP) (see Figure 1 (b)). The AP comprises a Squeeze Excitation Block (SEB) and a Spatial Attention Block (SAB). SEB applies channel attention to suppress irrelevant features and alleviate accumulated anatomical misalignment, while SAB further highlights salient spatial regions based on the refined feature maps, thereby reducing misalignment accumulation. In parallel, RP employs two convolutional layers for channel transformation and two sequential residual blocks to extract fine-grained anatomical details, supporting comprehensive and precise structural modeling (see Section 3.1). By adopting this design, FARM achieves high-level registration performance compared to a vanilla residual decoder without substantially increasing the number of learnable parameters. Moreover, to adaptively determine the number of optimization iterations for diverse image pairs, we propose the dual-stage Threshold-Controlled Iterative (TCI) strategy, which employs a sliding window of size $t$ to dynamically track the estimated deformation fields and quantifies each by the similarity difference between the warped moving image $I_{m}^{w}$ and the fixed image $I_{f}$. TCI then assesses the stability of deformation (i.e., first stage) by computing the standard deviation $\varepsilon_{l}$ of all similarity differences within this window. Upon reaching stability (see Section 3.2), TCI further evaluates convergence (i.e., second stage) by measuring the change $\Delta s$ between similarity differences of the two most recently estimated deformation fields within this window. TCI is designed to be efficient to enable adaptive determination of the number of iterations for each scale of the underlying image pair, without significantly increasing registration time (see Section 4.1). We coin the network that integrates FARM and TCI as the iterative Pyramid Estimation with Attention and Residuals (iPEAR) network. To the best of our knowledge, iPEAR is the first pyramid network to jointly model channel and spatial attention alongside residual blocks in its decoder, while dynamically determining the number of optimization iterations.
@@
I) We propose a dual-pathway FARM for decoding, comprising AP and RP, where AP suppresses irrelevant information and RP extracts fine-grained anatomical details.
II) We propose the dual-stage TCI strategy, which monitors the stability and convergence of recent registration results to adaptively determine the number of optimization iterations for image pairs with varying deformation requirements.
III) To assess the performance of iPEAR, we conduct experiments on three public brain Magnetic Resonance Imaging (MRI) datasets spanning different deformation scales and one public abdomen Computed Tomography (CT) dataset. The experimental results demonstrate that iPEAR outperforms SOTA registration models in accuracy across all datasets, while achieving on-par inference speed and model parameter size. Moreover, generalization and ablation studies further validate the effectiveness of FARM and TCI.
@@ 2 Related Work
Figure 2: Overall iPEAR model architecture. A weight-sharing encoder first extracts multi-scale feature maps $\{F_{l}\}$ and $\{M_{l}\}$ for the fixed image $I_{f}$ and the moving image $I_{m}$, respectively. In the subsequent decoding layers, iPEAR employs FARM to emphasize informative features and estimate the deformation field $\varphi_{l}$, while TCI adaptively determines the number of iterations for each decoding layer, enabling progressive optimization of $\varphi_{l}$. Notably, the encoder processes higher-resolution feature maps first ($F_{1}\rightarrow F_{4},M_{1}\rightarrow M_{4}$), while the decoder begins deformation estimation on the lowest‑resolution maps and proceeds to finer scales ($F_{4}\rightarrow F_{1},M_{4}\rightarrow M_{1}$).
In this section, we review the relevant literature.
@@
Deep Learning (DL) Registration Models: DL models have become the mainstream for solving registration tasks due to their advantages in inference speed and accuracy compared to conventional methods (Balakrishnan et al., 2019). Pioneer DL models primarily employed U-Net architectures, which estimate only a single deformation field at the final output layer and consequently struggle to accommodate complex deformations (Wang and Wang, 2025). To overcome this limitation, certain studies (De Vos et al., 2019; Zhao et al., 2019a, b) proposed cascaded registration networks, which employ multiple sub-networks to extract features at distinct resolutions and estimate the corresponding deformation fields. Although these models achieve moderate accuracy improvements, they perform repeated encoding and decoding, leading to redundant feature extraction and elevated computing overhead (Ma et al., 2024). To minimize redundant feature processing, recent studies (Liu et al., 2022) proposed the pyramid network, which adopts a single encoder followed by multiple decoding layers to estimate hierarchical deformation fields for continuous coarse‑to‑fine warping images. Although this architecture facilitates effective multi-scale feature extraction and progressive deformation estimation, it may exacerbate the propagation and accumulation of anatomical structure misalignments, especially those stemmed from irrelevant information in deep, coarse‑scale features, ultimately affecting the overall registration performance (Zhou et al., 2023). To better overcome this limitation, we propose FARM to suppress irrelevant features in the decoding layers, thereby mitigating misalignment propagation. Moreover, FARM excels at capturing anatomical details, ensuring the accuracy and consistency of deformation fields at each scale (see Section 3.1).
@@
Iterative Strategy in Pyramid Models: For small deformations, prior pyramid models (Kang et al., 2022) estimated only a single deformation field per decoding layer (each corresponding to one scale). However, under large deformations, a single estimation per scale is often insufficient to produce an accurate deformation field. Consequently, each scale requires iterative estimations to refine the deformation field. For example, Wang et al. (2024) proposed RDP, which performs a predefined number of iterations for each scale to refine the deformation field. Inevitably, this fixed number of iterations lacks adaptability to varying deformation requirements. To adaptively determine the number of iterations, Ma et al. (2024) proposed a single-stage mechanism, which evaluates convergence by monitoring differences between two successive registration results. However, relying on a single-metric check only captures partial information and may lead to premature iteration termination, potentially compromising the overall registration accuracy (see Tables 4 and F). To dynamically determine when to stop iterative optimization, we propose a dual-stage mechanism, which evaluates the stability of historical registration results and convergence subsequently (see Section 3.2 for more details).
3 iPEAR Dynamics
@@
The architecture of iPEAR is schematically illustrated in Figure 2. In this section, we first introduce FARM, which extracts anatomical structure details and learns to suppress irrelevant features to mitigate misalignment. We then introduce the dual-stage TCI strategy, which adaptively determines the number of iterations for optimization.
@@
Overview of iPEAR: iPEAR comprises a multi-scale encoder and a multi-layer decoder. iPEAR first employs a weight-sharing feature encoder to extract multi-scale feature map sets $\{F_{l}\}$ and $\{M_{l}\}$, where $l\in\{1,...,4\}$ (following prior arts in this field: (Kang et al., 2022; Ma et al., 2024)), from the fixed image $I_{f}$ and moving image $I_{m}$, respectively. Specifically, iPEAR adopts the encoder used in (Wang and Wang, 2025), which comprises four hierarchical convolutional encoder blocks. Each block includes a 3D convolutional layer followed by Neighborhood Attention (Hassani et al., 2023) (except for the first block), and then applies average pooling to halve the spatial dimensions of the feature maps. Then, for the $l$th scale (i.e., $l$th decoding layer), iPEAR applies FARM to process $\{F_{l}\}$ and $\{M_{l}\}$, extracting anatomical structure details and learning to suppress irrelevant features, to estimate the corresponding deformation fields (see Section 3.1). To adaptively determine the number of iterations in each decoding layer, iPEAR relies on TCI, which monitors the similarity difference of the recently estimated registration results (see Section 3.2).
Figure 3: The proposed FARM adopts a dual-pathway architecture. The AP is composed of SEB and SAB. SEB applies channel attention to suppress irrelevant features and alleviate accumulated anatomical misalignment. Building on these refined feature maps, SAB emphasizes salient spatial regions, which further reduces misalignment accumulation. In parallel, RP includes two convolutional layers for channel transformation and two sequential residual blocks that extract fine-grained anatomical details, supporting comprehensive and precise structural modeling.
3.1 Fused Attention-Residual Module (FARM)
Conventional residual convolutional structures, while effective at preserving anatomical consistency and modeling local details, typically apply uniform weighting across feature channels. This limits the network’s ability to emphasize subtle yet critical deformation cues that are essential for high-accuracy registration (Ma et al., 2024). In contrast, attention mechanisms excel at adaptively highlighting discriminative features and suppressing redundant information. However, when used alone, they often lack sufficient capability for robust local structure modeling and stable deformation estimation (Chen et al., 2022). Therefore, an effective decoder should combine the local modeling strength of residual convolutions with the adaptive feature selection of attention mechanisms to achieve high-level registration performance.
To this end, we propose FARM as the core component of the decoder for estimating deformation fields, which follows a dual-pathway design consisting of an Attention Pathway (AP) and a Residual Pathway (RP), as shown in Figure 3. Specifically, when $l=4$ and $k=1$, where $k$ denotes the current iteration number at the $l$th layer, FARM directly takes $F_{4}$ and $M_{4}$ as inputs to estimate the deformation field $\varphi_{4}^{1}\in\mathbb{R}^{C_{\varphi}\times H\times W\times D}$ via its dual-pathway architecture, where $C_{\varphi}$, $H$, $W$, and $D$ denote the number of channels, height, width, and depth of the deformation field, respectively. When $l\in\{1,2,3\}$ and $k=1$, FARM adopts the deformation field $\varphi_{l+1}$ of the $(l+1)$th layer to warp $M_{l}$ to obtain $M_{l}^{w}$ ($M_{l}^{w}=M_{l}\circ\varphi_{l+1}$), where $\circ$ denotes the deformation operation, implemented via the Spatial Transformer Network (Jaderberg et al., 2015). Then, FARM utilizes $F_{l}$ and $M_{l}^{w}$ to estimate the deformation field $\varphi_{l}^{1}\in\mathbb{R}^{C_{\varphi}\times H\times W\times D}$ for the $l$th layer through its dual pathways. Otherwise, FARM employs ${\varphi}_{l}^{k-1}$ of the (k1)(k-1)th iteration to warp $M_{l}$ to obtain $M_{l}^{w}$ ($M_{l}^{w}=M_{l}\circ\varphi_{l}^{k-1}$). Formally,
$\varphi_{l}^{k}=\text{FARM}(R_{l}),$ (1)
$R_{l}=\small{\left\{\begin{array}[]{ll}\left(|M_{4},F_{4}|\right),&l=4\land k=1,\\ \left(|M_{l}^{w},F_{l}|\right),&\text{Otherwise},\\ \end{array}\right.}$ (2)
where $|\cdot|$ denotes the concatenation operator, and the final expression of $\varphi_{l}^{k}$ is given later in (13).
AP consists of a Squeeze-and-Excitation Block (SEB) and a Spatial Attention Block (SAB), while RP comprises two convolutional layers for channel transformation and two sequential residual blocks. Specifically, AP extends the 2D SEB (Hu et al., 2018) to a 3D variant to suppress irrelevant features in $R_{l}$, thereby mitigating propagation and accumulation of anatomical misalignment. The 3D SEB first employs average pooling to compress spatial dimensions and extract global informative features, producing the channel descriptor vector $Z_{l}\in\mathbb{R}^{2C}$ as follows:
where $C$ denotes the number of channels of the image and $c\in\{1,\cdots,2C\}$ . Then, SEB applies two fully connected layers with LeakyReLU $\gamma$ and sigmoid $\sigma$ activation functions to compute the channel weight vector $S_{l}$ as follows:
$S_{l}=\sigma({W}_{2}\gamma({W}_{1}Z_{l})),$ (4)
where ${W}_{1}\in\mathbb{R}^{\frac{2C}{r}\times 2C}$ and ${W}_{2}\in\mathbb{R}^{2C\times\frac{2C}{r}}$. The hyperparameter $r$ denotes the reduction ratio, which is used to gauge the computational cost. The first linear layer reduces the dimensionality of $Z_{l}$ to $\frac{2C}{r}$, for efficient channel-wise compression and information aggregation. The second linear layer then restores the dimensionality back to $2C$. By adopting this design, SEB can model inter-channel dependencies and recalibrate channel responses. Consequently, SEB adaptively emphasizes deformation-sensitive channels while suppressing less informative ones, mitigating the tendency of uniform channel weighting to overlook subtle yet critical deformation cues (see Table 5). Subsequently, the resulting weight vector $S_{l}$ is applied to $R_{l}$ via per‑channel element‑wise multiplication $\ast$, yielding the enhanced feature map $O_{l}\in\mathbb{R}^{2C\times H\times W\times D}$ as follows:
$O_{l,c}=R_{l,c}\ast S_{l,c},c\in\{1,\cdots,2C\}.$ (5)
To further alleviate misalignment accumulation, SAB applies average pooling and max pooling to the enhanced feature map $O_{l}$ to obtain descriptor vectors $A_{l}$ and $M_{l}$, respectively. These two descriptors are then fused through a convolutional layer to produce the hybrid vector $H_{l}$, which captures complementary spatial cues. Formally,
$H_{l}=Conv(|A_{l},M_{l}|),$ (6)
$A_{l,c}=\frac{1}{H\cdot W\cdot D}\sum\nolimits_{i=1}^{H}\sum\nolimits_{j=1}^{W}\sum\nolimits_{d=1}^{D}O_{l,c,i,j,d},$ (7)
$M_{l,c}=\max_{i=\{1,\ldots,H\},j=\{1,\ldots,W\},d=\{1,\ldots,D\}}O_{l,c,i,j,d}.$ (8)
Subsequently, the descriptor vector $H_{l}$ is applied to $O_{l}$ via per‑channel element‑wise multiplication $\ast$, yielding the hybrid feature map $O^{{}^{\prime}}_{l}\in\mathbb{R}^{2C\times H\times W\times D}$ as follows:
$O{{}^{\prime}}_{l,c}=O_{l,c}\ast H_{l,c},c\in\{1,\cdots,2C\}.$ (9)
RP comprises a 3D convolutional layer ${Cl}_{1}$, two residual blocks ($Res_{1}$ and $Res_{2}$), and a final 3D convolutional layer $Cl_{2}$. Specifically, $R_{l}$ is first fed into a 3D convolutional layer, which adjusts the channel dimension from $2C$ to $C$, followed by a LeakyReLU function $\gamma$. The resulting features then pass through $Res_{1}$ and $Res_{2}$, each consisting of a 3D convolutional layer followed by $\gamma$. Finally, $Cl_{2}$ outputs the feature map $F_{\text{out}}$ with $\frac{C}{2}$ channels. Formally,
$Res_{i}(x)=\gamma(Conv(x)+x),i\in\{1,2\},$ (10)
$F_{o}=Cl_{2}^{C\rightarrow\frac{C}{2}}(Res_{2}(Res_{1}(\gamma(Cl_{1}^{2C\rightarrow C}(R_{l}))))),$ (11)
where $x$ denotes the input to the residual block. After obtaining $O^{{}^{\prime}}_{l}$ and $F_{o}$, FARM finally applies a convolutional layer $Cl_{3}$ to combine them and estimate the intermediate deformation field $\overline{\varphi}_{l}^{k}$ for the $k$th iteration as follows:
$\overline{\varphi}_{l}^{k}={Cl_{3}^{(\frac{C}{2}+2C\rightarrow 3)}}(|O^{{}^{\prime}}_{l},F_{o}|).$ (12)
After obtaining $\overline{\varphi}_{l}^{k}$, FARM forms the final $\varphi_{l}^{k}$ for the $k$th iteration by composing it with the deformation field $\varphi_{l}^{k-1}$ of the preceding iteration when $k\neq 1$. For the first iteration (i.e., $k=1$) with $l\neq 4$, FARM uses the final deformation field from the $(l+1)$th layer $\varphi_{l+1}$, in place of $\varphi_{l}^{k-1}$. When $l=4$ and $k=1$, FARM simply sets $\varphi_{l}^{k}=\overline{\varphi}_{l}^{k}$. Formally,
$\varphi_{l}^{k}=\left\{\begin{array}[]{ll}\varphi_{l}^{k-1}\circ\overline{\varphi}_{l}^{k}+\overline{\varphi}_{l}^{k},&k\neq 1,\\ \varphi_{l+1}\circ\overline{\varphi}_{l}^{k}+\overline{\varphi}_{l}^{k},&k=1,l\neq 4,\\ \overline{\varphi}_{l}^{k},&\text{Otherwise.}\end{array}\right$. (13)
Finally, when TCI terminates at the $k$th iteration, FARM upsamples $\varphi_{l}^{k}$ with a factor of 2 to obtain the final deformation $\varphi_{l}$ for the $l$th layer. Notably, unlike the decoder adopted by IIRP (Ma et al., 2024), which relies solely on multiple sequential residual blocks to estimate the deformation field, FARM integrates channel and spatial attention in parallel with a small number of residual blocks, designed to achieve enhanced performance through richer features (see Table 1). In addition, the results further validate that AP and RP effectively suppress irrelevant features and capture fine-grained anatomical details, respectively (see Appendix B for visualizations of feature attention patterns in AP and RP).
3.2 Threshold-Controlled Iterative (TCI) Strategy